Find the graph of the function f(n)=1-n from integers to integer


1
Expert's answer
2021-03-16T08:23:31-0400

Given a mapping "{\displaystyle f:X\to Y}", the graph of the mapping is the set


"{\displaystyle G(f)=\{(x,f(x))\mid x\in X\}} \subset {\displaystyle X\times Y}".


Let us find the graph of the function "f:\mathbb Z\to\mathbb Z", "f(n)=1-n":


"G(f)=\{ (n,f(n)) \mid n\in \mathbb Z \} =\{ (n, 1-n) \mid n\in \mathbb Z \}."


Let us sketch the garph: